OpenStudy (anonymous):

A point at the end of a spinning helicopter blade has a centripetal acceleration of 105.8 m/s2. If the velocity of the blade at this point is 23 m/s, how long is the helicopter blade?

OpenStudy (anonymous):

a = v^2/r r = v^2 / a square the speed, divide it by centripetal acceleration to get radius. You should multiply it by 2 to get the full length of the helicopter blade